# eTraM: Event-based Traffic Monitoring Dataset

### v1

@@ 4.1 Baseline Evaluation
To assess the performance of detectors of eTraM, we train them on of data and evaluate them on of validation and of test data. RVT and RED were trained from scratch over on A100, while YOLOv8 reached convergence in . We conducted separate evaluations to provide insights into how each model performs in different scenes and lighting conditions. This allows us to understand how these models handle diverse and changing environmental contexts.
Several key observations emerged from the evaluation results of tensor-based models shown in Table 2. Notably, the performance of vehicle detection consistently outperforms that of pedestrians across all scenes and models. The performance in the micro-mobility category exhibits significant variance, likely due to the smaller number of instances and the broad range of subjects captured within this category. During daytime, both vehicle and pedestrian detection results are consistent in intersections and roadways. However, in local streets, the performance of vehicle detection improves compared to other scenes due to fewer instances and reduced occlusion. Conversely, we observe a decline in the performance of pedestrian detection due to higher pedestrian densities leading to more inter-class occlusions.
During nighttime, the performance of vehicle detection in local streets and roadways remains consistent. Interestingly, pedestrian detection results improve significantly on local streets at night. This can be attributed to factors such as reduced pedestrian-to-pedestrian occlusion, instances being closer to the camera, and additional visual features like shadows that become more prominent at night due to street lights. Due to noise from various light sources, a drop in performance is observed at nighttime intersections. Despite this, the performance during nighttime is at par with daytime scenarios, with an increase in performance on vehicle detection on roadways at night. This increase could be due to reduced inter-class occlusion as we observe fewer vehicles in nighttime conditions.
Furthermore, tensor-based methods such as RED and RVT, which incorporate temporal information through Recurrent Neural Networks in their architecture, consistently outperform the conventional frame-based method of YOLOv8. Figure 4 shows an example of detection on eTraM. The evaluation highlights the challenges and strengths of various traffic monitoring scenarios and categories, particularly showcasing the effectiveness of event-based models in nighttime conditions.
@@ 4.2.1 Generalization on Night time
In [27], qualitative assessments were conducted to compare the generalization capabilities of event-based models against frame-based models in nighttime scenarios. We aim to quantitatively assess how well an event-based model trained on daytime data can perform in nighttime conditions. We conducted a controlled experiment where we trained two models for each architecture, one on of only daytime data and another with of daytime along with of nighttime data. The evaluation of both models was conducted on previously unseen nighttime data.
@@
The summarized results can be found in Table 3. We observe a consistent trend across every object class and architecture: models trained on data with nighttime sequences consistently outperform models trained solely on daytime sequences. Despite the expectation that event cameras would exhibit proficient performance in nighttime scenarios, these observations reveal that relying solely on daytime event data may not be sufficient. These models fail to attain performance levels comparable to those achieved by models trained on nighttime data with the model. This suggests that the unique challenges posed by nighttime conditions necessitate explicit training with relevant data to ensure optimal model performance. Upon closer inspection, we find that nighttime data has environmental interferences and distinct variations of noise. This could be inherent in nighttime data due to the heightened sensitivity of event cameras, which is a factor in this discrepancy.
@@ 4.2.2 Generalization on Unseen Scenes
As depicted in Table 4, we observe a comparable performance of the models across both the major categories considered. On evaluating the model on the “held out" test set, representing an entirely new and unseen traffic scene, the model’s generalization capability is evident. These values are at par with the performance on the “held in" test set. This similarity in results highlights the model’s ability to seamlessly extend its learning to previously unseen traffic scenes, validating its transfer capability across changing environments. This generalization to unseen intersections is pivotal for real-world deployment.
@@ 5.1 Event Cameras in Traffic Monitoring
Traditional cameras take continuous snapshots at a fixed frequency, potentially capturing individuals’ identities and sensitive information. In comparison, event cameras register events with no color and texture information, significantly reducing the probability of gathering sensitive information of any individual [27, 5]. Further, the advantages of event cameras extend to their performance in low light conditions and their robustness towards mitigating motion blur. As we demonstrate through our baseline evaluations in Table 2, event cameras display equally superior performance in nighttime conditions while maintaining a high temporal resolution that aids it in substantially reducing motion blur [31]. An essential requirement of traffic monitoring is the need for resource-efficient sensors. Since event cameras are designed to operate on a sparse data stream generated by significant visual changes, they exhibit lower memory requirements and power consumption compared to traditional frame cameras [2, 34]. This makes event cameras sustainable and cost-effective for continuous monitoring over extended periods.
@@ 5.2 Static and Ego Event-based Datasets
Next, we discuss the difference between static and ego-motion event-based datasets, emphasizing the significance of the former for traffic monitoring. In static datasets events are more likely to be associated with relevant traffic participants rather than extraneous background changes seen in ego-motion data, enhancing the visibility of these participants. This enhanced visibility is illustrated in Figure 6, which provides a qualitative comparison between events captured from a static and an ego-motion event camera. As shown, dynamic traffic participants from the static camera are better visible than those from the ego-motion camera. Another distinction is the field of view provided by static roadside datasets. Positioned at an elevated height, these cameras capture a broader scene, including far-away objects, while ego-motion datasets are constrained to the vehicle’s immediate surroundings. The elevated perspective of static datasets proves advantageous, enabling the detection of traffic participants at a distance and providing a comprehensive view of the traffic environment, as seen in Figure 7. This would be beneficial to a broad spectrum of applications, like traffic scene estimation [23, 38, 21], road safety monitoring [35] and traffic signal control [10].

### v2

# scale=1, color=gray, angle=0, opacity=0.5, position=current page.north, vshift=-1cm, contents= This paper has been accepted for publication at the IEEE Conference on Computer Vision and Pattern Recognition (CVPR), Seattle, 2024. ©IEEE eTraM: Event-based Traffic Monitoring Dataset

@@ 4.1 Baseline Evaluation
To assess the performance of detection methods on eTraM, we train them on of data and evaluate them on of validation and of test data. RVT and RED were trained from scratch over on A100, while YOLOv8 was trained for . We conducted separate evaluations to provide insights into how each model performs in different scenes and lighting conditions. This allows us to understand how these models handle diverse and changing environmental contexts.
Several key observations emerged from the evaluation results of tensor-based models shown in Table 2. Notably, the performance of vehicle detection consistently outperforms that of pedestrians across all scenes and models. The performance in the micro-mobility category exhibits significant variance, likely due to the smaller number of instances and the broad range of subjects captured within this category. During the daytime, both vehicle and pedestrian detection results are at par in intersections and roadways. In local streets, the performance of vehicle detection improves compared to other scenes due to fewer instances and reduced occlusion. Conversely, we observe a decline in the performance of pedestrian detection due to higher pedestrian densities leading to more inter-class occlusions.
During nighttime, the performance of vehicle detection in local streets and roadways remains consistent. Interestingly, pedestrian detection results improve significantly on local streets at night. This can be attributed to factors such as reduced pedestrian-to-pedestrian occlusion, instances being closer to the camera, and additional visual features like shadows that become more prominent at night due to street lights. Due to noise from various light sources, a drop in performance is observed at intersections during nighttime. Despite this, the performance during nighttime is at par with daytime scenarios, with an increase in performance on vehicle detection on roadways at night. This increase could be due to reduced inter-class occlusion as we observe fewer vehicles in nighttime conditions.
Furthermore, tensor-based methods such as RED and RVT, which incorporate temporal information through Recurrent Neural Networks in their architecture, consistently outperform the conventional frame-based method of YOLOv8. Figure 4 shows an example of detection on eTraM. The baseline results highlight the challenges and strengths of various traffic monitoring scenarios and categories, particularly showcasing the effectiveness of event-based models in nighttime conditions.
@@ 4.2.1 Generalization on Night time
In [28], qualitative assessments were conducted to compare the generalization capabilities of event-based models against frame-based models in nighttime scenarios. We aim to quantitatively assess how well event-based models RVT and RED, trained on daytime data, can perform in nighttime conditions. We conducted a controlled experiment where we trained two models for each architecture, one on of only daytime data and another with of daytime along with of nighttime data. Both models were evaluated on previously unseen nighttime data.
@@
The summarized results can be found in Table 3. We observe a consistent trend across every object class and architecture where models trained on data with nighttime sequences consistently outperform models trained solely on daytime sequences. Despite the expectation that event cameras would exhibit proficient performance in nighttime scenarios, these observations reveal that relying solely on daytime event data may not be sufficient. These models fail to attain performance levels comparable to those achieved by models trained on nighttime data with the model. This suggests that the unique challenges posed by nighttime conditions necessitate explicit training with relevant data to ensure optimal model performance. Upon closer inspection, we find that nighttime data has environmental interferences and distinct variations of noise. This could be inherent in nighttime data due to the heightened sensitivity of event cameras, which is a factor in this discrepancy.
@@ 4.2.2 Generalization on Unseen Scenes
As depicted in Table 4, we observe a comparable performance of the models across both major categories. On evaluating the model on the “held out" test set, representing an entirely new and unseen traffic scene, the model’s generalization capability is evident. These values are at par with the performance on the “held in" test set. This similarity in results highlights the model’s ability to seamlessly extend its learning to previously unseen traffic scenes, validating its transferability across changing environments. This generalization to unseen intersections is pivotal for real-world deployment.
@@ 5.1 Event Cameras in Traffic Monitoring
Traditional cameras take continuous snapshots at a fixed frequency, potentially capturing individuals’ identities and sensitive information. In comparison, event cameras register events with no color and texture information, significantly reducing the probability of gathering sensitive information of any individual [28, 5, 9]. Further, the advantages of event cameras extend to their performance in low light conditions and their robustness towards mitigating motion blur. As we demonstrate through our baseline evaluations in Table 2, event cameras display equally superior performance in nighttime conditions while maintaining a high temporal resolution that aids it in substantially reducing motion blur [32]. An essential requirement of traffic monitoring is the need for resource-efficient sensors. Since event cameras are designed to operate on a sparse data stream generated by significant visual changes, they exhibit lower memory requirements and power consumption compared to traditional frame cameras [2, 35]. This makes event cameras sustainable and cost-effective for continuous monitoring over extended periods.
@@ 5.2 Static and Ego Event-based Datasets
Next, we discuss the difference between static and ego-motion event-based datasets, emphasizing the significance of the former for traffic monitoring. In static datasets events are more likely to be associated with relevant traffic participants rather than extraneous background changes seen in ego-motion data, enhancing the visibility of these participants. This enhanced visibility is illustrated in Figure 6, which provides a qualitative comparison between events captured from a static and an ego-motion event camera. As shown, dynamic traffic participants from the static camera are better visible than those from the ego-motion camera. Another distinction is the field of view provided by static roadside datasets. Positioned at an elevated height, these cameras capture a broader scene and include far-away objects, while ego-motion datasets are constrained to the vehicle’s immediate surroundings. The elevated perspective of static datasets proves advantageous, enabling the detection of traffic participants at a distance and providing a comprehensive view of the traffic environment, as seen in Figure 7. This would be beneficial to a broad spectrum of applications, like traffic scene estimation [24, 39, 22], road safety monitoring [36] and traffic signal control [11].